# Video Soundtrack Generation by Aligning Emotions and Temporal Boundaries

### v1

@@ 3.2 Scene detection
During video-based inference, we first extract the video’s temporal boundaries, specifically its scene cut locations, using the FFmpeg software [^1] https://www.ffmpeg.org/. We then apply a difference filter to the extracted scene cuts and remove those occurring less than 4 seconds apart. The remaining timestamps serve as the input boundaries for the music generator.
3.3 Emotion matching
We use the VEMOCLAP emotion classifier to extract emotions from input videos as probability distributions over discrete emotion categories Sulun et al. (2024a). The music generator is conditioned on emotions represented as valence and arousal values ranging from -1 to 1, rather than discrete categories Sulun et al. (2022). Since the VEMOCLAP emotion classifier outputs discrete emotion categories, we first map them to the valence-arousal plane Russell (1980) before feeding them into our music generator.
Russell and Mehrabian conducted user studies to find the corresponding valence and arousal values of discrete emotion categories Russell and Mehrabian (1977). They presented their findings in a table that contains the mean and standard deviation values of valence and arousal for each categorical emotion. We extracted and used the values corresponding to Ekman’s six basic emotions, which are presented in Table 1 Ekman (1971).
@@ 3.3 Emotion matching
Table 1: Means and standard deviations of valence and arousal values corresponding to Ekman’s categorical emotions Ekman (1971), obtained from the user studies of Russell and Mehrabian Russell and Mehrabian (1977).
Using the output probabilities of the video classifier along with the means and standard deviations of each emotion, we construct a mixture of Gaussian distributions. Users can either sample from the mixture—first selecting an emotion category based on the output probabilities and then sampling from the corresponding Gaussian distribution—or use the mean of the mixture, which is computed as the weighted average of the emotion category means, with weights determined by the output probabilities. Additionally, we introduce a parameter that represents the maximum absolute value of means across all emotion categories. Using this parameter, we compute a single scaling coefficient applied to all valence-arousal distributions. This approach allows users to adjust the conditioning of the music generator, choosing between a wider or narrower range of emotions.
Group 1 Group 2
MR MQ EM TM OM MR MQ EM TM OM RT
Video2Music 2.521 1.854 2.125 1.958 2.167 2.708 1.771 1.854 1.979 2.083 1.607
CMT 1.854 2.438 2.000 2.125 2.083 1.792 2.562 2.458 2.312 2.354 3.554
EMSYNC (Ours) 1.625 1.708 1.875 1.917 1.750 1.500 1.667 1.688 1.708 1.562 1.417
Table 2: Group 1 consists of participants with a self-reported understanding of music theory, while Group 2 includes those without it. We use the following abbreviations: MR (music richness), MQ (music quality), EM (emotion match), TM (timing match), OM (overall match), and RT (runtime). Runtime measures the minutes required to generate music for a one-minute video. All other metrics represent average rankings, where 1 is the best possible and 3 is the worst possible ranking. Across all metrics, lower is better.
@@ 4 Experimental setup
As discussed in Section 2.3.1, conducting an objective evaluation is challenging when using unmatched video and MIDI datasets. We therefore focus on the subjective evaluation of our overall model. We conduct a user study to compare our results with two open-source state-of-the-art models that generate MIDI from arbitrary videos: Video2Music Kang et al. (2024) and CMT Di et al. (2021).
4.1 Dataset
We train our music generator on the Lakh Pianoroll-5 dataset Dong et al. (2018) with valence-arousal labels Sulun et al. (2022). For evaluation, we perform inference on videos from the Pittsburgh Advertisements Dataset (Ads) Hussain et al. (2017). We select this dataset because advertisements often rely on music to maximize viewer engagement Zander (2006). Additionally, since none of the compared methods have used this dataset, it ensures a fair comparison. Finally, the Ads dataset includes sentiment labels, allowing us to construct a well-balanced test set in terms of emotion.
We first filter the dataset by selecting videos associated with the four basic emotions commonly used in music emotion classification: cheerful, calm, angry, and sad Yang et al. (2008). We then remove videos shorter than 1 minute. To ensure an unbiased selection, we use YouTube IDs, which are generated randomly by YouTube. We sort these IDs alphabetically and select the first six videos from each emotion category, resulting in a total of 24 test videos. Finally, we trim all videos to a uniform duration of 1 minute.
@@ 4.2 Implementation details
We use the Music Transformer model with relative global attention in our music generator Huang et al. (2018). The model consists of 11 layers, 8 attention heads, and a dimensionality of 512. We train it with a context length of 1216 and a batch size of 64. Training is performed using cross-entropy loss with a learning rate of 2e-4 for the first 300k steps, followed by 5e-5 for the next 300k steps. We use the Adam optimizer with gradient clipping to a norm of 1 Kingma and Ba (2015). We do not apply any regularization methods and did not observe overfitting, as the model is trained on large-scale, densely labeled data where it predicts each token of its input sequence Sulun (2018).
For data augmentation, we transpose the pitches of all instruments, except drums, by a randomly chosen integer between -3 and 3, inclusive. We set our model size to be comparable to the baseline methods, with our model containing 37M parameters, compared to 39M in CMT Di et al. (2021) and 33M in Video2Music Kang et al. (2024). We set the maximum absolute value of means across all emotion categories to 0.8. For all models, we synthesize MIDI files into audio waveforms using Fluidsynth [^2] https://www.fluidsynth.com and apply peak audio normalization up to -3 dB.
@@ 4.3 Subjective evaluation
Using the 24 test videos, we generate accompanying music with the three models and create 12 survey pages, each containing two videos. For each video, the three music versions, generated by the compared models, are presented side by side with anonymized model names, with each model’s output appearing equally in the left, center, and right positions. We enroll 36 participants: 12 with self-reported knowledge of music theory, forming Group 1, and 24 without, forming Group 2. Each survey page and each video are viewed by one participant from Group 1 and two participants from Group 2.
We ask participants to rank the three models based on the standard criteria used in previous works Di et al. (2021); Kang et al. (2024): Music richness (MR): The richness and diversity of the music, independent of the video content. Music quality (MQ): The overall quality of the music, independent of the video content. Emotion match (EM): How well the music matches the video in terms of emotion. Timing match (TM): How well the music synchronizes with the video in terms of rhythm and timing. Overall match (OM): How well the music matches the video as a whole. Each model is assigned a unique ranking of 1 (best), 2, or 3 (worst), ensuring no ties. Finally, we report runtimes (RT) as the average time (in minutes) required to generate music for a one-minute video, including model initialization.
@@ 5 Results and conclusion
Table 2 compares EMSYNC, Video2Music Kang et al. (2024), and CMT Di et al. (2021) across all subjective criteria based on average rankings. A lower score represents a better ranking, with 1 being the best and 3 the worst possible.
EMSYNC consistently outperforms both baselines across all metrics for both Group 1 (music theory-aware) and Group 2. Group 2 rated EMSYNC even higher, particularly in music richness and emotion match, indicating that its strengths are even more apparent to general listeners. EMSYNC is also the fastest, with a runtime of 1.417 minutes, compared to 1.607 minutes for Video2Music and 3.554 minutes for CMT.
These results confirm that EMSYNC produces the most diverse and high-quality music with the best emotional alignment, rhythmic synchronization, and overall video compatibility while also being the most computationally efficient model.
By automatically generating soundtracks for user-provided videos, our work has the potential to streamline video production and enhance viewer engagement while offering a valuable framework for both machine learning researchers and multimedia content creators.

### v2

@@ 3.2 Scene detection
During video-based inference, we first extract the video’s temporal boundaries, specifically its scene cut locations. As scene detection is a solved problem, we employ the off-the-shelf FFmpeg software [^1] https://www.ffmpeg.org/. As we generate musical chords near scene boundaries, we aim to avoid an excessive number of densely spaced chords. We therefore apply a difference filter to the extracted scene cuts and remove those occurring less than 4 s apart. The remaining timestamps serve as the input boundaries for the music generator.
3.3 Emotion mapping
We use a pretrained emotion classifier to extract emotions from input videos as probability distributions over discrete emotion categories (Sulun, Viana, and Davies 2024). Our music generator is conditioned on emotions represented as valence and arousal values ranging from -1 to 1, rather than discrete categories. Since the emotion classifier outputs discrete emotion categories, we first map them to the valence-arousal plane (Russell 1980) before feeding them into the music generator.
Russell and Mehrabian conducted user studies to find the corresponding valence and arousal values of discrete emotion categories (1977). They presented their findings in a table that contains the mean and standard deviation values of valence and arousal for each categorical emotion. We extracted and used the values corresponding to Ekman’s six basic emotions (1971), which are presented in Table 1.
@@ 3.3 Emotion mapping
Table 1: Means and standard deviations of valence and arousal values corresponding to Ekman’s categorical emotions (1971), obtained from the user studies of Russell and Mehrabian (1977).
Using the output probabilities of the video classifier along with the means and standard deviations of each emotion, we construct a mixture of Gaussian distributions. Valence-arousal values can be obtained either by sampling an emotion category based on the output probabilities and then drawing from its corresponding Gaussian, or by using the mixture mean computed as the weighted average of the category means. Additionally, we introduce a parameter that represents the maximum absolute value of means across all emotion categories. Using this parameter, we compute a single scaling coefficient applied to all valence-arousal distributions. This approach allows users to adjust the conditioning of the music generator, choosing between a wider or narrower range of emotions.
Although not used in our work, it is possible to invert this mapping, that is, to map valence-arousal values to emotion categories by comparing the valence-arousal coordinates to each category’s distribution defined by its mean and standard deviation. This involves calculating a distance metric, such as Euclidean or Mahalanobis distance, or computing the likelihood assuming Gaussian distributions, and then assigning the category with the smallest distance or highest likelihood.
@@ 4 Experimental setup
We perform objective and subjective evaluations of our overall model using two datasets, against the open-source state-of-the-art models that generate MIDI from arbitrary videos: Video2Music (Kang, Poria, and Herremans 2024) and CMT (Di et al. 2021).
@@ 4.1 Implementation details
Our music generator consists of 11 layers, 8 attention heads, and a model dimensionality of 512. To ensure a fair comparison, we align our model size with the compared models: our model has 37M parameters, compared to 39M in CMT (Di et al. 2021) and 33M in Video2Music (Kang, Poria, and Herremans 2024). Training is done with a context length of 1216 and a batch size of 64. The model is implemented in PyTorch (Paszke et al. 2019) and trained on a single NVIDIA A100 80GB GPU.
For video-based inference in evaluation, we set the maximum absolute value of the emotion category means to 0.8. To ensure consistency, all methods generate MIDI outputs that are synthesized into audio waveforms using Fluidsynth [^2] https://www.fluidsynth.com, followed by peak normalization to -3 dB. The audio is then merged with the input video by overwriting its audio channel. For all methods, outputs are generated in a single run across all datasets using a fixed random seed.
We train our model using cross-entropy loss with a learning rate of 2e-4 for the first 300k steps, followed by 5e-5 for the next 300k steps. We use the Adam optimizer with gradient clipping to a norm of 1 (Kingma and Ba 2015). For data augmentation, we transpose the pitches of all instruments, except drums, by a randomly chosen number of semitones between -3 and 3, inclusive. We reserve 5% of the training data for validation. Hyperparameters are selected via grid search to minimize validation loss.
4.2 Evaluation datasets
For objective and subjective evaluation, we first use the EmoMV-C dataset, which consists of music videos containing music audio tracks to be used as ground truth for objective evaluation (Thao, Roig, and Herremans 2023). We exclude auditory features during the emotion classification to prevent data leakage. To assess the model’s generalization across different video types, we also include the Pittsburgh Advertisements (Ads) dataset, which contains advertisement videos that may or may not feature music (Hussain et al. 2017). As advertisements often rely on music to maximize viewer engagement, the Ads dataset provides a real-world scenario for evaluation. (Zander 2006).
We use the full EmoMV-C validation split, which contains 48 thirty-second videos. We filter the Ads dataset to match the total duration of EmoMV-C. We select videos associated with the four basic emotions commonly used in music emotion classification: cheerful, calm, angry, and sad (Yang et al. 2008). We then remove videos shorter than 1 minute. To ensure an unbiased selection, we use YouTube IDs, which are generated randomly by YouTube. We sort these IDs alphabetically and select the first six videos from each emotion category, resulting in 24 evaluation videos. Finally, we trim the videos to a uniform duration of 1 minute. The resulting evaluation videos are then fed into the compared models.
4.3 Objective evaluation
We first evaluate music similarity using the ground-truth audio from the EmoMV-C dataset. Although the ground truth is not in MIDI, all methods generate MIDI outputs that are synthesized to audio, enabling a fair comparison in the audio domain. We evaluate the Kullback-Leibler Divergence (KL) between the labels of generated and ground-truth music, calculated using a pretrained music tagging model (Won et al. 2020). This measures how closely the distribution of semantic tags in the generated music matches that of the ground-truth. We avoid using Fréchet Audio Distance, as it compares overall audio distributions suited for unconditional generation (Kilgour et al. 2019), whereas video-based music generation requires per-sample evaluation to assess alignment with each video’s unique content.
We additionally evaluate temporal audio-video alignment (AV-Align) score, which is the Intersection-over-Union between video and audio peaks (Yariv et al. 2024), using both EmoMV-C and Ads datasets. Video peaks are detected from the mean optical flow magnitude (Horn and Schunck 1981), while audio peaks correspond to audio onsets (Böck and Widmer 2013). We report AV-Align scores using windows of one video frame (33.4 ms at 30 FPS) and one second. Each window is applied to detect local peaks in both video and audio and then used to match peaks within the same temporal range for alignment.
@@ 4.4 Subjective evaluation
Using the evaluation videos from the EmoMV-C and the Ads datasets, we generate accompanying music with the three models and create survey pages, each containing two videos. For each video, the three music versions generated by the compared models are presented side by side with anonymized model names, with each model’s output appearing equally in the left, center, and right positions. We also provide output samples from eight videos online, with model names deanonymized and each model’s output appearing in the same position for clarity. [^3] https://em-sync.github.io
We enrolled 153 participants and asked them to rank the three models based on the standard criteria used in previous works (Di et al. 2021; Kang, Poria, and Herremans 2024): Music Richness (MR) — the richness and diversity of the music, independent of the video content; Music Quality (MQ) — the overall quality of the music, independent of the video content; Emotion Match (EM) — how well the music matches the video in terms of emotion; Timing Match (TM) — how well the music synchronizes with the video in terms of rhythm and timing; and Overall Match (OM) — how well the music matches the video as a whole. Each model is assigned a unique ranking of 1 (best), 2, or 3 (worst).
@@ 5 Results and conclusion
Table 2 presents objective evaluation results for EMSYNC, Video2Music (Kang, Poria, and Herremans 2024), and CMT (Di et al. 2021), on the EmoMV-C and Ads datasets. EMSYNC outperforms both baselines across a majority of the metrics. On EmoMV-C, it achieves the best frame-level alignment and the lowest KL divergence, indicating superior synchronization to video and similarity to ground-truth music. While Video2Music performs best on second-level alignment, EMSYNC remains competitive. On the Ads dataset, EMSYNC obtains the best scores across both alignment metrics, further demonstrating its robustness across different video domains. We don’t calculate KL divergence because the Ads dataset doesn’t include ground truth music.
Table 3 demonstrates the subjective results in terms of average user rankings with 95% confidence margins. A lower average score represents a better ranking, with 1 being the best and 3 the worst possible. EMSYNC achieves the best rankings across all metrics and datasets. The difference is more pronounced in video-music correspondence metrics such as emotion match and timing match, highlighting our model’s stronger alignment to video. For overall match, our method outperforms the others with non-overlapping 95% confidence intervals, reflecting a decisive user preference for EMSYNC in video-based music generation.
These results highlight EMSYNC’s effectiveness in generating music that aligns with the input video both temporally and emotionally, across objective and subjective evaluations. By automatically generating soundtracks for user-provided videos, our work has the potential to streamline video production and enhance viewer engagement while offering a valuable framework for both machine learning researchers and multimedia content creators.
Dataset EmoMV-C Ads
Metric AVf AVs KL AVf AVs
CMT 0.014 0.150 4.728 0.008 0.099
Video2Music 0.010 0.245 9.789 0.002 0.107
EMSYNC (Ours) 0.017 0.233 4.175 0.011 0.150
Table 2: Objective results: AVf and AVs denote audio-video alignment in frame and second-levels respectively, where higher is better. KL denotes KL-divergence where lower is better.
D M CMT Video2Music
EMSYNC
EmoMV-C
MR 1.870 ± 0.102 2.317 ± 0.108 1.812 ± 0.112
MQ 2.226 ± 0.099 1.952 ± 0.116 1.822 ± 0.113
EM 2.111 ± 0.100 2.043 ± 0.115 1.846 ± 0.117
TM 2.058 ± 0.098 2.067 ± 0.115 1.875 ± 0.120
OM 2.106 ± 0.100 2.087 ± 0.114 1.808 ± 0.116
Ads
MR 1.949 ± 0.153 2.439 ± 0.155 1.612 ± 0.140
MQ 2.469 ± 0.150 1.888 ± 0.144 1.643 ± 0.153
EM 2.214 ± 0.153 2.051 ± 0.171 1.735 ± 0.154
TM 2.245 ± 0.168 2.010 ± 0.159 1.745 ± 0.150
OM 2.255 ± 0.153 2.122 ± 0.166 1.622 ± 0.145
Table 3: Subjective results: Mean ranks ± 95% confidence margins for the subjective metrics, where 1 is the best possible and 3 is the worst possible ranking. The first two columns represent the dataset and metric, indicated by the headers D and M. The table is transposed for compactness.